OpenStudy (anonymous):

The intensity (I) of light received at a source varies inversely as the square of the distance (d) from the source. If the light intensity is 98 foot-candles at 27 feet, find the light intensity at 9 feet.

OpenStudy (saifoo.khan):

\[I = \frac{k}{d^2}\] \[98 = \frac{k}{27^2}\] k = 71442

OpenStudy (saifoo.khan):

\[I = \frac{71442}{9}\] I = 7938

OpenStudy (anonymous):

The correct answer was 882?? Oh man I hate these problems!!

OpenStudy (dumbcow):

saifoo is correct in the set up just forgot to square the 9 71442/81 = 882

OpenStudy (saifoo.khan):

Ohhh! SORRYY!